4FeS2 + 11O2 > 2Fe2O3 + 8SO2

Given 75.0 grams of iron (IV) sulfide, how many grams of sulfur dioxide are produced?
1
Expert's answer
2020-03-02T06:57:45-0500

"\nu=m/M;"

"\nu(FeS2) =75/120=0.625 mol."

0.625 : 4="\nu(SO2)" : 8

"\nu(SO2) =0.625*8/4=1.25 mol."

"m=\nu*M=1.25*64=80 g."


Answer: 80 g.